OpenStudy (anonymous):

HELP! I need help with these absolute problems. Can anyone help me? Very appreciated.

OpenStudy (anonymous):

m<=-13 or m>=3

OpenStudy (anonymous):

Could you help me with a few more?

OpenStudy (kewlgeek555):

Well, if |m+5| has to be greater than 8, than m should equal 3.1 and above.

OpenStudy (anonymous):

there`s an equality sign too @kewlgeek555

OpenStudy (anonymous):

ya tell me @Brookebrookiee

OpenStudy (anonymous):

@divu.mkr Okay, I need help with a few more.. |m+5|>=8 |3m-6|=12 |2x-3|<=7 21=|2x|+3 |2y-5|>=9 2|x-7|-6=6 t-(2-t)|>10